# Spectrum of Schrödinger operators on subcovering graphs

### v1

@@ Abstract.
We consider discrete Schrödinger operators with periodic potentials on periodic graphs. Their spectra consist of a finite number of bands. By ”rolling up” a periodic graph along some appropriate directions we obtain periodic graphs of smaller dimensions called subcovering graphs. For example, rolling up a planar hexagonal lattice along different directions will lead to nanotubes with various chiralities. We show that the subcovering graph is asymptotically isospectral to the original periodic graph as the length of the ”chiral” (roll up) vectors tends to infinity and get asymptotics of the band edges of the Schrödinger operator on the subcovering graph. We also obtain a criterion for the subcovering graph to be just isospectral to the original periodic graph. By isospectrality of periodic graphs we mean that the spectra of the Schrödinger operators on the graphs consist of the same number of bands and the corresponding bands coincide as sets.
@@ 1. Introduction
In this paper we consider discrete Schrödinger operators with periodic potentials on arbitrary periodic graphs. Their spectra consist of a finite number of segments called bands. The $n$-th band is the range of a band function $\lambda_{n}(k)$ describing the dependence of electron energy on the multidimensional parameter $k$ called quasimomentum which takes its values in the Brillouin zone $(-\pi,\pi]^{d}$, where $d$ is the dimension of the periodic graph (i.e., the rank of its period lattice). The multi-valued function $k\mapsto\{\lambda_{n}(k)\}$ is called the dispersion relation. By ”rolling up” a periodic graph in some appropriate directions we obtain periodic graphs of smaller dimensions called subcovering graphs. The aim of this paper is to describe connections between spectra of the Schrödinger operators on a periodic graph and its subcoverings.
@@
We describe our main results.
$\bullet$ We formulate a criterion for the subcovering graph to be isospectral to the original periodic graph. By isospectrality of periodic graphs we mean that the spectra of the Schrödinger operators on the graphs consist of the same number of bands and the corresponding bands coincide as sets.
$\bullet$ We show that subcovering graphs are asymptotically isospectral to their original periodic graphs as the length of the chiral (roll up) vectors tends to infinity, and obtain asymptotics of the band edges for the Schrödinger operator on the subcovering graph.
The proof is based on the restriction procedure applied to the dispersion relation for the Schrödinger operator on the original periodic graph which was used in [KP07] for the hexagonal lattice.
@@
Note that, due to the band structure of the spectrum of the Schrödinger operators on periodic graphs, there are some notions of isospectrality for periodic graphs: Floquet isospectrality, Fermi isospectrality, isospectrality for the periodic problem (see Section 2.5 for the definitions and references).
@@
i) It is known (see, e.g., Theorem 2.1.ii in [KS19]) that the first spectral band $\sigma_{1}(H)=[\lambda_{1}^{-},\lambda_{1}^{+}]$ is never a flat band and its lower point $\lambda_{1}^{-}=\lambda_{1}(0)$.
ii) The band functions $\lambda_{j}(\cdot)$, $j\in{\mathbb{N}}_{\nu}$, are even, i.e.,
$\lambda_{j}(-k)=\lambda_{j}(k),\qquad\forall\,k\in{\mathcal{B}},$
see, e.g., Lemma 2.9.i in [KKR17].
@@ 1.3. Spectra of Schrödinger operators on subcovering graphs
The coordinate vectors $t_{s}\in{\mathbb{Z}}^{d}$ of $\mathfrak{t}_{s}\in\Gamma$ with respect to the basis $\mathfrak{a}_{1},\ldots,\mathfrak{a}_{d}$ of the lattice $\Gamma$ will be called the chiral indices of the subcovering ${\mathcal{G}}_{\mathfrak{t}}$, and the matrix ${\mathcal{T}}$ whose rows are the chiral indices $t_{s}$, $s\in{\mathbb{N}}_{d_{o}}$, will be called the chiral matrix of ${\mathcal{G}}_{\mathfrak{t}}$.
@@
$\bullet$ we formulate a criterion for the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ to be not only asymptotically isospectral but just isospectral to the original periodic graph ${\mathcal{G}}$ (Proposition 2.4 and Corollary 2.5);
@@
We illustrate the obtained results by some examples of periodic graphs and their subcoverings with asymptotically close spectra. In particular, we describe all subcoverings of the $d$-dimensional lattice, the hexagonal lattice and the diamond lattice which are isospectral to the original periodic graphs (Examples 2.8, 2.10 and 2.12, respectively).
@@ 2.1. Primitive sets in a lattice
2.2. Asymptotically isospectral subcoverings
@@
Let ${\mathcal{G}}$ be a $\Gamma$-periodic graph with the fundamental graph ${\mathcal{G}}_{*}={\mathcal{G}}/\Gamma$, where $\Gamma$ is a lattice of rank $d$. From now on we will consider only subcoverings ${\mathcal{G}}_{\mathfrak{t}}$ of ${\mathcal{G}}$ with primitive sets $\mathfrak{t}\subset\Gamma$ of chiral vectors $\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}$, $d_{o}<d$. Any primitive set $\mathfrak{t}=\{\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\}$ of $\Gamma$ can be extended to a basis
$\{\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}},\mathfrak{t}_{d_{o}+1},\ldots,\mathfrak{t}_{d}\}$
of the lattice $\Gamma$. Let
$\bullet$ $\Gamma_{\mathfrak{t}}$ be the sublattice of $\Gamma$ with the basis $\{\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\}$, and
$\bullet$ $\Gamma_{o}$ be the sublattice of $\Gamma$ with the basis $\{\mathfrak{t}_{d_{o}+1},\ldots,\mathfrak{t}_{d}\}$.
Then the subcovering ${\mathcal{G}}_{\mathfrak{t}}={\mathcal{G}}/\Gamma_{\mathfrak{t}}$ is a $\Gamma_{o}$-periodic graph with the fundamental graph
${\mathcal{G}}_{\mathfrak{t}}/\Gamma_{o}=({\mathcal{G}}/\Gamma_{\mathfrak{t}})/\Gamma_{o}={\mathcal{G}}/\Gamma={\mathcal{G}}_{*}.$ (2.1)
Thus, the periodic graph ${\mathcal{G}}$ and any its subcovering ${\mathcal{G}}_{\mathfrak{t}}$ with a primitive set $\mathfrak{t}$ have the same fundamental graph ${\mathcal{G}}_{*}=({\mathcal{V}}_{*},{\mathcal{E}}_{*})$, and, consequently, the spectra of the Schrödinger operators $H=\Delta+Q$ on ${\mathcal{G}}$ and $H_{\mathfrak{t}}=\Delta_{\mathfrak{t}}+Q$ on ${\mathcal{G}}_{\mathfrak{t}}$ consist of the same number $\nu:=\#{\mathcal{V}}_{*}$ of bands.
Recall that the chiral matrix ${\mathcal{T}}$ of the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ is defined by (1.8), (1.9). Denote by $\tau$ the smallest eigenvalue of $({\mathcal{T}}{\mathcal{T}}^{\top})^{1/2}$, where ${\mathcal{T}}^{\top}$ is the transpose of ${\mathcal{T}}$. Note that, since $\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}$ are linearly independent, the matrix ${\mathcal{T}}{\mathcal{T}}^{\top}$ is positive definite.
We show that the Schrödinger operators $H$ and $H_{\mathfrak{t}}$ have asymptotically close spectra as $\tau\to\infty$.
Proposition 2.2.
$\begin{array}[]{l}\displaystyle\sigma(H_{\mathfrak{t}})=\bigcup_{j=1}^{\nu}\sigma_{j}(H_{\mathfrak{t}}),\qquad\sigma_{j}(H_{\mathfrak{t}})=\lambda_{j}({\mathcal{B}}_{\mathfrak{t}})=\big{[}\lambda_{j}^{-}(\mathfrak{t}),\lambda_{j}^{+}(\mathfrak{t})\big{]},\qquad\nu=\#{\mathcal{V}}_{*},\\ \displaystyle\lambda_{j}^{-}(\mathfrak{t})=\min_{k\in{\mathcal{B}}_{\mathfrak{t}}}\lambda_{j}(k)\geqslant\min_{k\in{\mathcal{B}}}\lambda_{j}(k)=:\lambda_{j}^{-},\\[10.0pt] \displaystyle\lambda_{j}^{+}(\mathfrak{t})=\max_{k\in{\mathcal{B}}_{\mathfrak{t}}}\lambda_{j}(k)\leqslant\max_{k\in{\mathcal{B}}}\lambda_{j}(k)=:\lambda_{j}^{+},\end{array}$ (2.2)
where $\lambda_{j}(\cdot)$ are the band functions of the Schrödinger operator $H=\Delta+Q$ on the periodic graph ${\mathcal{G}}$, and ${\mathcal{B}}_{\mathfrak{t}}\subset{\mathcal{B}}=(-\pi,\pi]^{d}$ is defined by (1.10). Moreover,
$\lim\limits_{\tau\to\infty}\lambda_{j}^{-}(\mathfrak{t})=\lambda_{j}^{-},\qquad\lim\limits_{\tau\to\infty}\lambda_{j}^{+}(\mathfrak{t})=\lambda_{j}^{+},\qquad\forall\,j\in{\mathbb{N}}_{\nu},$ (2.3)
where $\tau$ is the smallest eigenvalue of $({\mathcal{T}}{\mathcal{T}}^{\top})^{1/2}$, and ${\mathcal{T}}$ is the chiral matrix of ${\mathcal{G}}_{\mathfrak{t}}$ defined by (1.8), (1.9).
Remark 2.3.
i) The condition $\tau\to\infty$ implies, in particular, that the Euclidean norm $|t_{s}|$ of each chiral index $t_{s}\in{\mathbb{Z}}^{d}$, see (1.8), is large enough. Indeed, the diagonal entries of the matrix ${\mathcal{T}}{\mathcal{T}}^{\top}$ are $|t_{1}|^{2},\ldots,|t_{d_{o}}|^{2}$. Then the smallest eigenvalue $\tau^{2}$ of ${\mathcal{T}}{\mathcal{T}}^{\top}$ satisfies
$\tau^{2}\leqslant\min\big{\{}|t_{1}|^{2},\ldots,|t_{d_{o}}|^{2}\big{\}},$
ii) Proposition 2.2 shows that the spectral bands
of the Schrödinger operators $H$ on a periodic graph ${\mathcal{G}}$ and $H_{\mathfrak{t}}$ on its subcovering ${\mathcal{G}}_{\mathfrak{t}}$ with a primitive set $\mathfrak{t}$ satisfy $\sigma_{j}(H_{\mathfrak{t}})\subseteq\sigma_{j}(H)$. Due to (2.3), the band edges $\lambda_{j}^{\pm}(\mathfrak{t})$ of $H_{\mathfrak{t}}$ are asymptotically close to the corresponding band edges $\lambda_{j}^{\pm}$ of $H$ as $\tau\to\infty$. Then we say that ${\mathcal{G}}_{\mathfrak{t}}$ is asymptotically isospectral to ${\mathcal{G}}$.
iii) If the set $\mathfrak{t}\subset\Gamma$ is not primitive, then Proposition 2.2 is not true any more. For example, if $d_{o}=1$ and the single chiral vector $\mathfrak{t}=p\mathfrak{t}_{o}$ for some primitive $\mathfrak{t}_{o}\in\Gamma$ and integer $p\geqslant 2$, then the fundamental graph of the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ has $p\nu$ vertices, and, consequently, the spectrum of the Schrödinger operator $H_{\mathfrak{t}}$ on ${\mathcal{G}}_{\mathfrak{t}}$ consists of $p\nu$ spectral bands, see Example 4.2. Note that ${\mathcal{G}}_{\mathfrak{t}_{o}}$ is a $p$-fold subcovering of ${\mathcal{G}}_{p\mathfrak{t}_{o}}$, which yields that $\sigma(H_{\mathfrak{t}_{o}})\subseteq\sigma(H_{p\mathfrak{t}_{o}})$, see, e.g., [A95].
@@
Now we formulate a simple criterion for the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ to be isospectral to the original periodic graph ${\mathcal{G}}$. Denote by $K_{j}^{\pm}$ the level sets corresponding to the band edges $\lambda_{j}^{\pm}$ of the Schrödinger operator $H$ on ${\mathcal{G}}$, i.e.,
@@
Let $\mathfrak{t}=\{\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\}$ be a primitive set of the lattice $\Gamma$, and $j\in{\mathbb{N}}_{\nu}$. Then the band edge $\lambda_{j}^{\pm}(\mathfrak{t})$ of the Schrödinger operator $H_{\mathfrak{t}}$ on the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ coincides with the corresponding band edge $\lambda_{j}^{\pm}$ of the Schrödinger operator $H$ on the original $\Gamma$-periodic graph ${\mathcal{G}}$, i.e., $\lambda_{j}^{\pm}(\mathfrak{t})=\lambda_{j}^{\pm}$, if and only if
@@
for some $k_{o}\in K_{j}^{\pm}$, where $K_{j}^{\pm}$ are given by (2.4) and ${\mathcal{T}}$ is the chiral matrix of ${\mathcal{G}}_{\mathfrak{t}}$ defined by (1.8), (1.9). In particular,
i) If $0\in K_{j}^{\pm}$, then $\lambda_{j}^{\pm}(\mathfrak{t})=\lambda_{j}^{\pm}$.
ii) If $(\pi,\ldots,\pi)\in K_{j}^{\pm}$ and the sum of the entries in each row of ${\mathcal{T}}$ is even, then $\lambda_{j}^{\pm}(\mathfrak{t})=\lambda_{j}^{\pm}$.
The following criterion of isospectrality is a direct consequence of Proposition 2.4.
Corollary 2.5.
Let $\mathfrak{t}=\{\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\}$ be a primitive set of the lattice $\Gamma$. Then a $\Gamma$-periodic graph ${\mathcal{G}}$ is isospectral to its subcovering ${\mathcal{G}}_{\mathfrak{t}}$, i.e., $\sigma(H)=\sigma(H_{\mathfrak{t}})$, if and only if for each $j\in{\mathbb{N}}_{\nu}$ the condition (2.5) holds for some $k^{-}\in K_{j}^{-}$ and for some $k^{+}\in K_{j}^{+}$, where $K_{j}^{\pm}$ are given by (2.4).
@@
Here $\mathrm{Hess}\,\lambda_{j}(k_{o})$ is the Hessian of $\lambda_{j}$ at $k_{o}$, ${\mathcal{T}}$ is the chiral matrix of ${\mathcal{G}}_{\mathfrak{t}}$ defined by (1.8), (1.9), and $\tau$ is the smallest eigenvalue of $({\mathcal{T}}{\mathcal{T}}^{\top})^{1/2}$; ||$\cdot$|| ⋅ | denotes the Euclidean norm of a vector.
@@
i) Note that the second term in the asymptotics (2.6) is equal to zero if and only if ${\mathcal{T}}k_{o}\in 2\pi{\mathbb{Z}}^{d_{o}}$. This agrees with Proposition 2.4 (see the condition (2.5)).
@@
iii) For a particular case when $d_{o}=1$ and the single chiral index has the specific form $t=(t_{1},\ldots,t_{d-1},-1)\in{\mathbb{Z}}^{d}$, i.e., $t_{d}=-1$, the statements of Proposition 2.4, Corollary 2.5 and Theorem 2.6 were obtained in [S24], see also Remark 1.7.
iv) Assumption A1 is often used to establish many important properties of periodic operators such as electron’s effective masses [AM76], Green’s function asymptotics [KKR17, K18] and other issues. Although in the continuous case it is commonly believed that the band edges are non-degenerate for ”generic” potentials, in discrete settings there are known counterexamples [FK18, BK20], as well as various positive results in some cases [DKS20].
@@
In this section we illustrate the obtained results by some simple examples of periodic graphs and their asymptotically isospectral and just isospectral subcoverings. The proofs of the examples are based on Proposition 2.4 and Theorem 2.6 and given in Section 4.
@@
see Lemma 4.1. We describe the spectrum of the Schrödinger operator $H_{\mathfrak{t}}=\Delta_{\mathfrak{t}}+Q$ on subcoverings (nanotubes) ${\bf G}_{\mathfrak{t}}$ of the hexagonal lattice ${\bf G}$.
@@
i) For any primitive $\mathfrak{t}\in\Gamma$, the nanotube ${\bf G}_{\mathfrak{t}}$ is asymptotically isospectral to the hexagonal lattice ${\bf G}$ as $|t|\to\infty$. If $t_{1}-t_{2}\in 3{\mathbb{Z}}$, then the Schrödinger operators $H$ on ${\bf G}$ and $H_{\mathfrak{t}}$ on ${\bf G}_{\mathfrak{t}}$ have the same spectrum, i.e., ${\bf G}$ and ${\bf G}_{\mathfrak{t}}$ are just isospectral. If $\mathfrak{t}$ is not primitive, i.e., $\mathfrak{t}=p\mathfrak{t}_{o}$ for some primitive $\mathfrak{t}_{o}\in\Gamma$ and integer $p\geqslant 2$, then the spectrum of the Schrödinger operator $H_{\mathfrak{t}}$ on ${\bf G}_{\mathfrak{t}}$ consists of $2p$ spectral bands, some of which may be degenerate (see also Remark 2.3.iii and Example 4.2).
@@
$\lambda_{1}^{+}(\mathfrak{t})=4d-\frac{\pi^{2}}{\tau^{2}}+O\big{(}\frac{1}{\tau^{4}}\big{)},\qquad\textrm{as}\qquad\tau=|\mathfrak{t}_{1}|\to\infty.$ (2.12)

### v2

@@ Abstract.
We consider discrete Schrödinger operators with real periodic potentials on periodic graphs. The spectra of the operators consist of a finite number of bands. By ”rolling up” a periodic graph along some appropriate directions we obtain periodic graphs of smaller dimensions called subcovering graphs. For example, rolling up a planar hexagonal lattice along different directions will lead to nanotubes with various chiralities. We describe connections between spectra of the Schrödinger operators on a periodic graph and its subcoverings. In particular, we provide a simple criterion for the subcovering graph to be isospectral to the original periodic graph. By isospectrality of periodic graphs we mean that the spectra of the Schrödinger operators on the graphs consist of the same number of bands and the corresponding bands coincide as sets. We also obtain asymptotics of the band edges of the Schrödinger operator on the subcovering graph as the ”chiral” (roll up) vectors are long enough.
@@ 1. Introduction
In this paper we consider discrete Schrödinger operators with real periodic potentials on arbitrary periodic graphs. Their spectra consist of a finite number of segments called bands. The $n$-th band is the range of a band function $\lambda_{n}(k)$ describing the dependence of electron energy on the multidimensional parameter $k$ called quasimomentum which takes its values in the Brillouin zone $(-\pi,\pi]^{d}$, where $d$ is the dimension of the periodic graph (i.e., the rank of its period lattice). The multi-valued function $k\mapsto\{\lambda_{n}(k)\}$ is called the dispersion relation. By ”rolling up” a periodic graph in some appropriate directions we obtain periodic graphs of smaller dimensions called subcovering graphs. The aim of this paper is to describe connections between spectra of the Schrödinger operators on a periodic graph and its subcoverings.
@@
In this paper we describe spectra of the Schrödinger operators on subcovering graphs using some apriori information about the spectrum of the operator on the original periodic graph. In particular, we formulate a criterion for the subcovering graph to be isospectral to the original periodic graph. By isospectrality of periodic graphs we mean that the spectra of the Schrödinger operators on the graphs consist of the same number of bands and the corresponding bands coincide as sets. We get asymptotics of the band edges of the Schrödinger operator on the subcoverings as the length of the chiral (roll up) vectors tends to infinity. The proof is based on the restriction procedure applied to the dispersion relation for the Schrödinger operator on the original periodic graph which was used in [KP07] for the hexagonal lattice, and on the construction of the minimum norm solution of a linear system of equations corresponding to this restriction.
@@
i) The first spectral band $\sigma_{1}(H)=[\lambda_{1}^{-},\lambda_{1}^{+}]$ is never a flat band and $\lambda_{1}^{-}=\lambda_{1}(0)$, see, e.g., Theorem 2.1.ii in [KS19].
ii) The band functions $\lambda_{j}(\cdot)$, $j\in{\mathbb{N}}_{\nu}$, are even (see Lemma 2.9.i in [KKR17]), i.e.,
$\lambda_{j}(-k)=\lambda_{j}(k),\qquad\forall\,k\in{\mathcal{B}}.$
@@ 1.3. Spectra of Schrödinger operators on subcovering graphs
The coordinate vectors $t_{s}\in{\mathbb{Z}}^{d}$ of the chiral vectors $\mathfrak{t}_{s}\in\Gamma$ with respect to the basis $\mathfrak{a}_{1},\ldots,\mathfrak{a}_{d}$ of the lattice $\Gamma$ will be called the chiral indices of the subcovering ${\mathcal{G}}_{\mathfrak{t}}$, and the matrix ${\mathcal{T}}$ whose rows are the chiral indices $t_{s}$, $s\in{\mathbb{N}}_{d_{o}}$, will be called the chiral matrix of ${\mathcal{G}}_{\mathfrak{t}}$.
@@
$\bullet$ we formulate a simple criterion for the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ to be isospectral to the original periodic graph ${\mathcal{G}}$ (Proposition 2.2.i);
@@
We illustrate the obtained results by some examples of periodic graphs and their subcoverings. In particular, we describe all subcoverings of the $d$-dimensional lattice, the hexagonal lattice and the diamond lattice which are isospectral to the original periodic graphs (Examples 2.6, 2.8 and 2.10, respectively).
@@ 2.1. Primitive sets in a lattice
2.2. Spectra of Schrödinger operators on subcoverings with primitive sets of chiral vectors
@@
Let ${\mathcal{G}}$ be a $\Gamma$-periodic graph with the fundamental graph ${\mathcal{G}}_{*}={\mathcal{G}}/\Gamma=({\mathcal{V}}_{*},{\mathcal{E}}_{*})$, where $\Gamma$ is a lattice of rank $d$. From now on we will consider only subcoverings ${\mathcal{G}}_{\mathfrak{t}}$ of ${\mathcal{G}}$ with primitive sets $\mathfrak{t}\subset\Gamma$ of chiral vectors $\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}$, $d_{o}<d$.
@@
We describe the spectrum of the Schrödinger operators $H_{\mathfrak{t}}$ on the subcovering ${\mathcal{G}}_{\mathfrak{t}}$. Denote by $K_{j}^{\pm}$ the level sets corresponding to the band edges $\lambda_{j}^{\pm}$ of the Schrödinger operator $H$ on ${\mathcal{G}}$, i.e.,
@@
$\sigma(H_{\mathfrak{t}})=\bigcup_{j=1}^{\nu}\sigma_{j}(H_{\mathfrak{t}}),\qquad\sigma_{j}(H_{\mathfrak{t}})=\big{[}\lambda_{j}^{-}(\mathfrak{t}),\lambda_{j}^{+}(\mathfrak{t})\big{]}\subseteq\big{[}\lambda_{j}^{-},\lambda_{j}^{+}\big{]},\qquad\nu=\#{\mathcal{V}}_{*},$ (2.2)
where $\lambda_{j}^{\pm}$ are the band edges of the Schrödinger operator $H=\Delta+Q$ on ${\mathcal{G}}$ defined in (1.6). Moreover, the following statements hold true:
i) $\lambda_{j}^{\pm}(\mathfrak{t})=\lambda_{j}^{\pm}$ if and only if
@@
for some $k_{o}\in K_{j}^{\pm}$, where $K_{j}^{\pm}$ are given by (2.1), and ${\mathcal{T}}$ is the chiral matrix of ${\mathcal{G}}_{\mathfrak{t}}$ defined by (1.8), (1.9). In particular,
$\bullet$ If $0\in K_{j}^{\pm}$, then $\lambda_{j}^{\pm}(\mathfrak{t})=\lambda_{j}^{\pm}$.
$\bullet$ If $(\pi,\ldots,\pi)\in K_{j}^{\pm}$ and the sum of the entries in each row of ${\mathcal{T}}$ is even, then $\lambda_{j}^{\pm}(\mathfrak{t})=\lambda_{j}^{\pm}$.
ii) Let $\tau$ be the smallest eigenvalue of $({\mathcal{T}}{\mathcal{T}}^{\top})^{1/2}$, where ${\mathcal{T}}^{\top}$ is the transpose of ${\mathcal{T}}$. Then
$\lim\limits_{\tau\to\infty}\lambda_{j}^{\pm}(\mathfrak{t})=\lambda_{j}^{\pm},\qquad\forall\,j\in{\mathbb{N}}_{\nu}.$ (2.4)
Remark 2.3.
i) Proposition 2.2 shows that the spectral bands
of the Schrödinger operators $H$ on a periodic graph ${\mathcal{G}}$ and $H_{\mathfrak{t}}$ on its subcovering ${\mathcal{G}}_{\mathfrak{t}}$ with a primitive set $\mathfrak{t}$ satisfy $\sigma_{j}(H_{\mathfrak{t}})\subseteq\sigma_{j}(H)$.
ii) From Proposition 2.2 it also follows that a $\Gamma$-periodic graph ${\mathcal{G}}$ is isospectral to its subcovering ${\mathcal{G}}_{\mathfrak{t}}$, if and only if for each $j\in{\mathbb{N}}_{\nu}$ the condition (2.3) holds for some $k^{-}\in K_{j}^{-}$ and for some $k^{+}\in K_{j}^{+}$, where $K_{j}^{\pm}$ are given by (2.1). By isospectrality of periodic graphs we mean that the spectra of the Schrödinger operators on the graphs consist of the same number of bands and the corresponding bands coincide as sets.
iii) If the set $\mathfrak{t}\subset\Gamma$ is not primitive, then Proposition 2.2 is not true any more. For example, if $d_{o}=1$ and the single chiral vector $\mathfrak{t}=p\mathfrak{t}_{o}$ for some primitive $\mathfrak{t}_{o}\in\Gamma$ and integer $p\geqslant 2$, then the fundamental graph of the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ has $p\nu$ vertices, and, consequently, the spectrum of the Schrödinger operator $H_{\mathfrak{t}}$ on ${\mathcal{G}}_{\mathfrak{t}}$ consists of $p\nu$ spectral bands. Note that ${\mathcal{G}}_{\mathfrak{t}_{o}}$ is a $p$-fold subcovering of ${\mathcal{G}}_{p\mathfrak{t}_{o}}$, which yields that $\sigma(H_{\mathfrak{t}_{o}})\subseteq\sigma(H_{p\mathfrak{t}_{o}})$, see, e.g., [A95].
iv) Since $\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}$ are linearly independent, the matrix ${\mathcal{T}}{\mathcal{T}}^{\top}$ is positive definite. The condition $\tau\to\infty$ implies, in particular, that the Euclidean norm $|t_{s}|$ of each chiral index $t_{s}\in{\mathbb{Z}}^{d}$, see (1.8), tends to infinity. Indeed, the diagonal entries of the matrix ${\mathcal{T}}{\mathcal{T}}^{\top}$ are $|t_{1}|^{2},\ldots,|t_{d_{o}}|^{2}$. Then the smallest eigenvalue $\tau^{2}$ of ${\mathcal{T}}{\mathcal{T}}^{\top}$ satisfies
$\tau^{2}\leqslant\min\big{\{}|t_{1}|^{2},\ldots,|t_{d_{o}}|^{2}\big{\}},$
v) Due to (2.4), the band edges $\lambda_{j}^{\pm}(\mathfrak{t})$ of $H_{\mathfrak{t}}$ are asymptotically close to the corresponding band edges $\lambda_{j}^{\pm}$ of $H$ as $\tau\to\infty$. Then we say that ${\mathcal{G}}_{\mathfrak{t}}$ is asymptotically isospectral to ${\mathcal{G}}$ as $\tau\to\infty$. Notion of asymptotic isospectrality for Schrödinger operators on periodic graphs was introduced in [S24] when studying the behavior of the spectrum of the Schrödinger operators on periodic graphs under perturbations of graphs by adding edges (between existing vertices) in a periodic way.
vi) Due to the band structure of the spectrum of the Schrödinger operators on periodic graphs, there are some notions of isospectrality for periodic graphs. The problem of Floquet isospectrality for the discrete Schrödinger operators on the lattice ${\mathbb{Z}}^{d}$ was studied in [GKT93, K89, L23]. Fermi isospectrality on ${\mathbb{Z}}^{d}$ was discussed in [L24].
@@
Here $\mathrm{Hess}\,\lambda_{j}(k_{o})$ is the Hessian of $\lambda_{j}$ at $k_{o}$; ${\mathcal{T}}$ is the chiral matrix of ${\mathcal{G}}_{\mathfrak{t}}$ defined by (1.8), (1.9); $\tau$ is the smallest eigenvalue of $({\mathcal{T}}{\mathcal{T}}^{\top})^{1/2}$, and ||$\cdot$|| ⋅ | denotes the Euclidean norm of a vector.
@@
i) The second term in the asymptotics (2.5) is equal to zero if and only if ${\mathcal{T}}k_{o}\in 2\pi{\mathbb{Z}}^{d_{o}}$. This agrees with Proposition 2.2, see the condition (2.3).
@@
iii) Although Assumption A1 is often used to establish many important properties of periodic operators (for example, the notion of electron’s effective masses in solid state physics [AM76]), for some periodic discrete graphs this assumption fails (see examples in [FK18, BK20]).
iv) For a particular case when $d_{o}=1$ and the single chiral index has the specific form $t=(t_{1},\ldots,t_{d-1},-1)\in{\mathbb{Z}}^{d}$, i.e., $t_{d}=-1$, the statements of Proposition 2.2 and Theorem 2.4 were obtained in [S24], see also Remark 1.7.
@@
In this section we illustrate the obtained results by some simple examples of periodic graphs and their subcoverings. The proofs of the examples are based on Proposition 2.2 and Theorem 2.4 and given in Section 4.
@@
see, e.g., Lemma 4.1 in [S24]. We describe the spectrum of the Schrödinger operator $H_{\mathfrak{t}}=\Delta_{\mathfrak{t}}+Q$ on subcoverings (nanotubes) ${\bf G}_{\mathfrak{t}}$ of the hexagonal lattice ${\bf G}$.
@@
i) For any primitive $\mathfrak{t}\in\Gamma$, the nanotube ${\bf G}_{\mathfrak{t}}$ is asymptotically isospectral to the hexagonal lattice ${\bf G}$ as $|t|\to\infty$. If $t_{1}-t_{2}\in 3{\mathbb{Z}}$, then the Schrödinger operators $H$ on ${\bf G}$ and $H_{\mathfrak{t}}$ on ${\bf G}_{\mathfrak{t}}$ have the same spectrum, i.e., ${\bf G}$ and ${\bf G}_{\mathfrak{t}}$ are just isospectral. If $\mathfrak{t}$ is not primitive, i.e., $\mathfrak{t}=p\mathfrak{t}_{o}$ for some primitive $\mathfrak{t}_{o}\in\Gamma$ and integer $p\geqslant 2$, then the spectrum of the Schrödinger operator $H_{\mathfrak{t}}$ on ${\bf G}_{\mathfrak{t}}$ consists of $2p$ spectral bands, some of which may be degenerate. But, due to (1.12), the inclusion $\sigma(H_{p\mathfrak{t}_{o}})\subseteq\sigma(H)$ still holds.
For example, if ${\bf G}_{\mathfrak{t}}$ is the zig-zag nanotube with the non-primitive vector $\mathfrak{t}=2\mathfrak{a}_{1}$ (see Fig. 1c), then the spectrum of the Schrödinger operator $H_{\mathfrak{t}}=\Delta_{\mathfrak{t}}+Q$ on ${\bf G}_{\mathfrak{t}}$ with the periodic potential $Q$ defined by (2.7) has the form
$\begin{array}[]{l}\sigma(H_{\mathfrak{t}})=\sigma_{ac}(H_{\mathfrak{t}})\cup\sigma_{fb}(H_{\mathfrak{t}}),\qquad\sigma_{fb}(H_{\mathfrak{t}})=\{3\pm\sqrt{1+q^{2}}\,\},\\[8.0pt] \sigma_{ac}(H_{\mathfrak{t}})=\big{[}3-\sqrt{9+q^{2}}\,,3-\sqrt{1+q^{2}}\,\big{]}\cup\big{[}3+\sqrt{1+q^{2}}\,,3+\sqrt{9+q^{2}}\,\big{]},\end{array}$
see, e.g., [KK10]. The spectrum $\sigma(H_{\mathfrak{t}})$ consists of four spectral bands (since the fundamental graph of ${\bf G}_{2\mathfrak{a}_{1}}$ has four vertices, see Fig. 1e), two of which are flat bands.
@@
$\lambda_{1}^{+}(\mathfrak{t})=4d-\dfrac{\pi^{2}}{\tau^{2}}+O\big{(}\frac{1}{\tau^{4}}\big{)},\qquad\textrm{as}\qquad\tau=|\mathfrak{t}_{1}|\to\infty.$ (2.11)